Sol Summary Report – February 23th

Crew 275 Sol Summary Report 23Feb2023

Sol: 11

Summary Title: The sweetness of success

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Today, the EVA Crew going to North Ridge had the opportunity to study a 3D map of the area before starting their exploration tasks. They reported that they felt almost like they already knew the place and managed to reach all the targeted checkpoints very quickly. This seems to show that such 3D maps generated by drones are very effective in training crews to explore a place they have never been to before. It was actually so effective that they came back about an hour earlier than planned. We decided to spend this spare time cooking, and ended up with a very tasty brioche and an amazing cake!

Inside the Hab, we continued our activities, including installing the software dedicated to AI-assisted echography on a more powerful tablet. It is now even easier for non-trained subjects like us to perform ultrasounds on each other.

In the context of our experiment with the Royal Institute of Technology (Sweden), our Crew Engineer is now working on a location tracking system that will enable the possibility to correlate our stress and emotions to our position in the station.

Look Ahead Plan:

If the weather conditions allow it, we would like to perform the tasks planned for EVA #9 (which was cancelled) tomorrow morning (Sol 12). We would thus re-install the antenna of Mega-Ares and the weather instruments that have been removed during EVA #7. As no rain is forecasted for the next few days, we will also re-install the LOAC.

On Saturday (Sol 13), we would like to perform our first EVA dedicated to the MetMet experiment. The plan would be to collect geological samples based on the information given by a user-friendly instrument measuring the magnetic susceptibility and conductivity of rocks (the aim being to check the interest of a rock sample before collecting it). In addition to this, we would change the batteries of the atmospheric instruments, because we do not plan any EVA for Sunday (Sol 14) and the batteries cannot last more than 48 hours in cold weather.

Anomalies in work: None.

Weather: Windy,low -6°C/22°F, high 3°C/37°F.

Crew Physical Status: Optimal.

EVA: EVA #10 at North Ridge went smoothly, and the EVA Crew managed to reach all the desired checkpoints very quickly.

Reports to be filed: Sol Summary, EVA Request, EVA Report, Operations Report, Journalist Report, GreenHab Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site